OpenStudy (anonymous):

I need some help with an AP Physics problem. I have a .025 kg block moving at .2 m/s and it hits a .01 kg block moving .15 m/s, I need to find the resulting velocity of each block. @Mathematics

OpenStudy (anonymous):

at what angle do they hit?

OpenStudy (anonymous):

This is a momentum question?

OpenStudy (anonymous):

Yes, this is a momentum question. They are in a straight line going in the same direction.

OpenStudy (anonymous):

This is the actual problem word for word if it helps. A 25.0-g object moving to the right at 20.0 cm/s overtakes and collides elastically with a 15.0-g object moving the in same direction at 15.0 cm/s. Find the velocity of each object after the collision.

OpenStudy (anonymous):

p = mv always. I think, p1 = 0.005 kg m/s p2 = 0.0015 kgm/s When p1 and p2 collide m1u1 + m2u2 = m1v1 + m2v2 (0.025)(0.2) + (0.01)(0.15) = (0.025)(v1) + (0.01)(v2) That is as far as I can go. There probably is some piece of information missing.

OpenStudy (anonymous):

But if the collision is perfectly inelastic, then v1 = v2 and you solve for v, v = [6.5 X 10^-3] / (0.025 + 0.01) = 0.185 m/s

OpenStudy (anonymous):

The answer 17.1 cm/s for the 25 g object and 22.1 cm/s for 10 g gram object. I'm going to try and work backwards and figure something out. Thank you for your effort.

OpenStudy (anonymous):

that is the same as 18.5 cm/s

OpenStudy (anonymous):

I'm going to try to work with the equations and substitute. Maybe something will work out.

OpenStudy (anonymous):

Very well. Good luck.

OpenStudy (anonymous):

Helpful hint... momentum AND kinetic energy are conserved. So maybe try and find a way to relate the two equations: \[p=mv\]\[KE=(1/2)mv{^2}\] Keeping in mind the 'initial' and 'final' should balance.

OpenStudy (stormfire1):

This may help you see the steps...problem #37 is almost exactly the same: http://physics.ucsd.edu/students/courses/summer2009/session1/managed/session1/physics1a/documents/Ch6_solutions.pdf